# Understanding Pre-training and Fine-tuning from Loss Landscape Perspective

### v1

@@ Abstract
Recent studies have revealed that the loss landscape of large language models resembles a basin, within which the models perform nearly identically, and outside of which they lose all their capabilities. In this work, we conduct further studies on the loss landscape of large language models. We discover that pre-training creates a "basic capability" basin, and subsequent fine-tuning creates "specific capability" basins (e.g., math, safety, coding) within the basic capability basin. We further investigate two types of loss landscapes: the most-case landscape (i.e., the landscape along most directions) and the worst-case landscape (i.e., the landscape along the worst direction). We argue that as long as benign fine-tuning remains within the most-case basin, it will not compromise previous capabilities. Similarly, any fine-tuning (including the adversarial one) that stays within the worst-case basin would not compromise previous capabilities. Finally, we theoretically demonstrate that the size of the most-case basin can bound the size of the worst-case basin and the robustness with respect to input perturbations. We also show that, due to the over-parameterization property of current large language models, one can easily enlarge the basins by five times.
@@ 1 Introduction
Large language models (LLMs) have garnered significant attention in recent years for their remarkable performance across diverse applications [52, 4, 23, 42]. These models typically undergo a pre-training phase with extensive datasets to acquire foundational knowledge, followed by multiple alignment stages using high-quality, domain-specific data to activate specialized capabilities [10, 52, 53]. In this work, we investigate the intriguing alignment brittleness phenomenon. In particular:
  • Why does fine-tuning with benign data sometimes compromise capabilities acquired through prior alignment [58, 22, 50, 9, 46, 31, 41, 44]?
  • Why does fine-tuning with adversarial data, even for just a few steps, rapidly destroy all capabilities of large language models [58, 59, 67, 33, 39, 32, 34, 16]?
  • Why are large language models easily jailbroken in white-box settings, and how does this relate to the above issues [77, 57, 14, 3]?
We propose that these questions may be partially explained by the loss landscape of large language models [40]. We examine two types of loss landscapes: the most-case landscape, which reflects capacity degradation when parameters move along most directions, and the worst-case landscape, which captures degradation along the most detrimental direction.
@@
Figure 1: The most-case loss landscape of different models. Specific benchmarks and visualization details are provided in Sec. 3.2. As shown, the loss landscape of LLMs resembles a basin, within which models perform nearly identically and outside of which they lose all capabilities. The basic capability basin, endowed during pre-training, is consistently the largest. Fine-tuning creates specific capability basins (e.g., math, safety, coding) within this basic capability basin.
@@
As shown in Fig. 1, the most-case landscape resembles a basin, within which models perform nearly identically and outside of which they rapidly lose all capabilities [54]. This aligns with recent findings that large language models can resist common noise perturbations [49, 29, 64]. The pre-training phase establishes a "basic capability basin" that endows the model with fundamental language comprehension and conversational abilities. Subsequent alignment stages sequentially create specific capability basins (e.g., coding, math, safety) within this basic capability basin. Based on this observation, we propose that as long as benign fine-tuning remains within a specific capability basin, the parameters remain in that basin, preserving prior capabilities.
@@
As illustrated in Fig. 2, the worst-case landscape is consistently sharp, such that even a small fine-tuning step can move parameters outside the basin, resulting in the loss of all capabilities. This phenomenon resembles prior explanations for adversarial examples: in high-dimensional spaces, there exists a direction that causes rapid degradation, despite most directions being safe [65, 27, 47]. The parameter dimensions of large language models are significantly larger than those of earlier smaller models, making the worst-case direction potentially more detrimental. Furthermore, we propose that a lack of robustness to worst-case parameter perturbations implies vulnerability to input perturbations, such as jailbreaking. Let $\bm{W}$ denote the embedding layers. Given that the embedding layers of current large language models are onto transformations (i.e., $\bm{W}$ is column full-rank) [11], if there exists a perturbation $\bm{\delta_{W}}$ such that the model with weights $\bm{W}+\bm{\delta_{W}}$ is not robust, then there always exists an input perturbation $\bm{\delta_{x}}$ such that the model with input $\bm{x}+\bm{\delta_{x}}$ is not robust, as $\bm{W}\bm{x}+\bm{\delta_{W}}\bm{x}$ and $\bm{W}\bm{x}+\bm{W}\bm{\delta_{x}}$ can yield the same vector [74]. This explains the vulnerability of large language models to both jailbreaking [77] and fine-tuning attacks [58].
@@
Through exploratory preliminary studies, we find that one can construct a smooth model such that the size of the worst-case basin is theoretically lower bounded by the size of the most-case basin. This implies that we can derive a theoretical upper bound on performance degradation for any fine-tuning, even along the worst direction, as well as for input jailbreaking. Combined with our conjecture that benign fine-tuning preserves capabilities within the most-case basin, we conclude that enlarging the most-case basin enhances benign fine-tuning, mitigates harmful fine-tuning, and improves robustness against input jailbreaking. Additionally, we demonstrate that the basins can be expanded five times with ease, likely due to the over-parameterization property of neural networks [8, 2]. We hope our work sheds light on the relationships among loss landscapes, robustness to benign and harmful fine-tuning, and jailbreaking, and that our proposed theoretical lower bounds and optimization strategies inspire future large-scale studies on pre-training and fine-tuning.
2 Alignment Brittleness of LLMs
@@
The alignment of LLMs plays a crucial role in ensuring adherence to safety and ethical standards in their applications [5, 68, 36]. During the early stages of LLM advancement, researchers developed various paradigms, like reinforcement learning [20, 6] or red-teaming [55, 48], to build their alignment, which was initially believed to be sufficient to solve the alignment problems. However, a series of recent discoveries revealed that the current alignment of LLMs is shallow and superficial [69, 58, 72, 77, 57, 75]. Though performing aligned values in regular deployments, their alignment ability may be easily broken during fine-tuning or reasoning phases, particularly in the following three aspects:
@@
Normal fine-tuning [^1] This is what previous work called “Benign Fine-tuning”. In this paper, benign fine-tuning refers to another type of fine-tuning defined in Sec. 3.4. Supervised Fine-Tuning (SFT) on task-specific datasets has become a common paradigm for applying pre-trained base LLMs in various practical scenarios. However, even after achieving desirable alignment after pre-training, these base models may significantly forget their alignment after fine-tuning. For example, the Llama-2-7b-Chat model raises their harmfulness response rate from 5.5% to 31.8% after fine-tuning on the Alpaca dataset with only one epoch [58].
@@
Adversarial fine-tuning. Furthermore, this superficial alignment can be easily undermined by a small amount of adversarial data. By exploiting the inherent vulnerabilities in the alignment of large language models (LLMs), malicious actors can use a minimal number of harmful examples, such as instructions that encourage harmful behavior or shift identities. For instance, fine-tuning GPT-3.5-turbo on a maliciously crafted dataset containing just 10 harmful examples for 5 epochs can raise the harmfulness rate from 1.8% to 88.8% [58], effectively dismantling the safety mechanisms of sufficiently aligned LLMs.
@@
Input-space jailbreaking. Finally, LLMs after fine-tuning still suffer from input-space attacks, which are known as jailbreaking attacks. For example, adversaries can utilize optimization-based methods to induce the model to answer harmful outputs, even in black-box settings [70, 12, 3, 16]. Similarly, adversaries can exploit other input modules like vision inputs in VLLMs to induce the model to generate any target harmful content [21, 73, 56], with an attack success rate of 100%.
@@
3 A Closer Look at the Loss Landscape of LLMs
In this work, we explore these unaddressed questions through the perspective of loss landscapes, which has achieved notable success in understanding various dynamics of conventional neural networks [40, 54].
3.1 Visualization of the Loss Landscape
The loss landscape visualizes the performance degradation of a neural network with respect to parameter perturbations. Formally, let $f_{\bm{\theta}}$ denote a language model with parameters $\bm{\theta}\in\mathbb{R}^{d}$, and let $J_{\mathcal{D}}$ represent the benchmark functional on a dataset $\mathcal{D}$, defined as $J_{\mathcal{D}}(f_{\bm{\theta}})=\mathbb{E}_{\bm{x}\in\mathcal{D}}[J(f_{\bm{\theta}}(\bm{x}))]$. This functional takes the language model $f_{\bm{\theta}}$ as input and returns a benchmark value on $\mathcal{D}$, characterizing its specific capabilities. The loss landscape is thus a visualization of the high-dimensional function $J_{\mathcal{D}}:\mathbb{R}^{d}\to\mathbb{R}$.
Benchmarks. To assess the diverse capabilities of a given model, we adopt the following benchmarks: MMLU [30] for basic language proficiency, GSM8K [17] for mathematical reasoning, HumanEval [24] for coding ability, and AdvBench [77] for safety performance. Since the values of different benchmarks are not directly comparable, we normalize each landscape to the interval $[0,1]$ and invert benchmarks where higher values indicate better performance, unifying them such that lower values indicate better performance for consistent visualization.
Models. We visualize the loss landscape of three cutting-edge models, including Llama-3.1-8B [23], Qwen-2.5-7B [71] and Mistral-8B-2410 [37]. We also analyze the relationship between loss landscape and model size in Sec. D.2.
However, since the function $J_{\mathcal{D}}(\bm{\theta})$ is high-dimensional, directly visualizing a $d$-dimensional landscape is computationally expensive and unintuitive [40]. To address this challenge, researchers typically visualize the loss landscape along a specific direction $\bm{\delta}\in\mathbb{R}^{d}$, reducing the problem to visualizing a single-variable function [40, 28, 35, 62]:
$L(\alpha)=J_{\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta}).$ (1)
In this work, we investigate two types of loss landscapes defined by the choice of direction ($\bm{\delta}$):
3.2 Most-case Loss Landscape
The most-case loss landscape uses a uniformly random direction $\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})$, visualizing the single-variable function:
$L(\alpha)=J_{\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta}),\;\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I}).$ (2)
Empirically, we observe that different directions $\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})$ yield nearly identical results. Since $\mathcal{N}(\bm{0},\bm{I})$ represents uniformly random directions, this suggests that most directions produce similar landscapes [^2] This can be validated through hypothesis testing; see the Sec. D.3 for details.. Therefore, we designate this as the most-case loss landscape.
General Geometry. As shown in Fig. 1, the most-case loss landscape for each capability resembles a basin, within which the models perform nearly identically, and outside of which they rapidly lose all capabilities [54]. The pre-training stage creates a "basic capability basin" that endows the model with fundamental language comprehension and conversational abilities. Subsequent alignment stages sequentially establish specific capability basins (e.g., coding [24], math [17], safety [77]) within this basic capability basin. Based on this observation, we argue that as long as subsequent benign fine-tuning remains within the basin of a specific capability, the parameters will remain within this basin and thus will not compromise those capabilities.
@@ 3.2 Most-case Loss Landscape
Figure 2: The worst-case loss landscape of the three models. As shown, moving even a small distance along the worst-case direction rapidly degrades all capabilities of LLMs. (Due to all curves reaching the maximum loss at the smallest scale, they completely overlap.)

### v2

# Unveiling the Basin-Like Loss Landscape in Large Language Models

@@ Abstract
We discover the emergence of basins in the loss landscape of large language models. As model scale increases, LLMs become progressively more resilient to random perturbations in the parameter space, giving rise to expansive stability regions where models exhibit nearly identical performance, but outside of which their capabilities collapse. We observe that pre-training creates a basic capability basin, and subsequent alignment fine-tuning forms specific capability basins (e.g., safety, math, coding). Thus, we argue that benign fine-tuning confined to the basin should preserve prior capabilities. Besides, we also analyze the loss landscape for worst-case directions, which is consistently sharp and detrimental. We find that adversarial fine-tuning moves along the nearly worst-case directions, thus rapidly degrading model capabilities. Finally, we provide a theoretical analysis demonstrating that the basin size bounds the performance degradation of any fine-tuning, including the adversarial ones, while also guaranteeing the model robustness w.r.t. input perturbations, suggesting the benefit of enlarging basins.
@@ 1 Introduction
Large Language Models (LLMs) have garnered significant attention in recent years for their remarkable performance across numerous applications (OpenAI, 2023; Anthropic, 2024; Dubey et al., 2024; Liu et al., 2024a). LLMs typically undergo a pre-training phase with extensive datasets to acquire foundational knowledge, followed by multiple alignment stages using high-quality, domain-specific data to activate specialized capabilities (Brown et al., 2020; OpenAI, 2023; Ouyang et al., 2022). In this work, we investigate the intriguing alignment brittleness phenomenon. In particular:
  • Why does fine-tuning with benign data sometimes compromise capabilities acquired during prior alignment (Qi et al., 2023; Du et al., 2024; Mukhoti et al., 2023; Bianchi et al., 2023; Lyu et al., 2024; Hsu et al., 2024; Li et al., 2025; Liu et al., 2024b)?
  • Why does fine-tuning with adversarial data, even for just a few steps, destroy all capabilities of LLMs (Qi et al., 2023; Rosati et al., 2024; Wang et al., 2024a; Huang et al., 2024b; Leong et al., 2024; Huang et al., 2024a; c; Wu et al., 2025)?
  • Why are LLMs easily jailbroken in white-box settings, and how does this relate to the above issues (Zou et al., 2023; Qi et al., 2024; Chen et al., 2025a; Andriushchenko et al., 2024)?
We posit that these issues can be partially explained by the loss landscape of LLMs (Li et al., 2018). To investigate this, we analyze two complementary perspectives: the most-case landscape, capturing capacity degradation when parameters move along most directions, and the worst-case landscape, reflecting degradation along the worst direction. Our main findings are summarized below.
@@
Figure 1: The most-case loss landscape of different models. Specific benchmarks and visualization details are provided in Sec.˜3.2. As shown, the loss landscape of LLMs resembles a basin, within which models perform nearly identically and outside of which they lose all capabilities.
@@
As shown in Fig.˜1, the most-case landscape exhibits a basin-like structure: models perform nearly identically within the basin, but rapidly lose capabilities once outside (Peng et al., 2024). This basin gradually emerges and expands as model size increases (see Fig.˜7), consistent with recent evidence that LLMs are robust to common parameter perturbations (Men et al., 2024; Gromov et al., 2024; Sun et al., 2023). In particular, pre-training forms a broad “basic capability basin” that confers fundamental language and conversational skills, while subsequent alignment stages carve out narrower “specific capability basins” (e.g., safety, math, coding) in proximity to it. These observations suggest that benign fine-tuning constrained within the basin preserves prior capabilities.
@@
As illustrated in Fig.˜2, the worst-case landscape is consistently sharp, such that even a small fine-tuning step can move parameters outside the basin, leading to the loss of all capabilities. This resembles prior explanations for adversarial examples: in high-dimensional spaces, with high probability there exists a direction that causes rapid degradation, despite most directions being safe (Daniely & Shacham, 2020; Bubeck et al., 2021). The parameter dimensions of LLMs are significantly larger than those of earlier smaller models, making the worst-case direction potentially more detrimental. Furthermore, we argue that a lack of robustness to worst-case parameter perturbations implies vulnerability to input perturbations, such as jailbreaking. Let $\bm{W}$ denote the embedding layers. Given that the embedding layers of current LLMs are onto transformations (i.e., $\bm{W}$ is column full-rank) (Carlini et al., 2024), if there exists a perturbation $\bm{\delta_{W}}$ such that the model with weights $\bm{W}+\bm{\delta_{W}}$ is not robust, there always exists an input perturbation $\bm{\delta_{x}}$ such that the model at input $\bm{x}+\bm{\delta_{x}}$ is not robust, as $\bm{W}\bm{x}+\bm{\delta_{W}}\bm{x}$ and $\bm{W}\bm{x}+\bm{W}\bm{\delta_{x}}$ can yield the same output (Zhang et al., 2024a). This explains LLM vulnerabilities to both jailbreaking (Zou et al., 2023) and fine-tuning attacks (Qi et al., 2023).
@@
Through above exploratory studies, we construct a smooth model such that the performance degradation along the worst-case direction is theoretically bounded by the size of the most-case basin. This implies that we can derive a theoretical bound on performance degradation for any fine-tuning and input jailbreaking. Together with our conjecture that benign fine-tuning preserves capabilities within the most-case basin, we conclude that enlarging the most-case basin ❶ enhances benign fine-tuning, ❷ mitigates adversarial fine-tuning, and ❸ improves robustness against input jailbreaking. Additionally, we demonstrate that the basins can be readily enlarged during pre-training, likely due to the over-parameterization property of neural networks (Belkin et al., 2019; Allen-Zhu et al., 2019). We hope that our work sheds light on the relationships between loss landscapes, robustness to benign and harmful fine-tuning, jailbreaking, and our proposed theoretical lower bounds and optimization strategies inspire future large-scale studies on pre-training and fine-tuning.
2 Background: Alignment Brittleness of LLMs
@@
The alignment of LLMs plays a crucial role in ensuring adherence to safety and ethical standards in their applications (Anwar et al., 2024; Wang et al., 2024b; Ji et al., 2023). During the early stages of LLM advancement, researchers developed various paradigms, like reinforcement learning (Dai et al., 2024; Bai et al., 2022) or red-teaming (Perez et al., 2022; Mehrabi et al., 2023), to build their alignment, which was initially believed to be sufficient to solve the alignment problems. However, a series of recent discoveries revealed that the current alignment of LLMs is shallow and superficial (Wei et al., 2023a; Qi et al., 2023; Yang et al., 2023; Zou et al., 2023; Qi et al., 2024; Zhang et al., 2024b). Although models exhibit aligned behavior in regular settings, their alignment can be easily compromised during fine-tuning or inference, particularly in the following three aspects.
@@
Normal fine-tuning [^1] This is what previous work called “benign fine-tuning”. In this paper, benign fine-tuning refers to another type of fine-tuning defined in Sec. 3.4. Supervised Fine-Tuning (SFT) on task-specific datasets has become a common paradigm for adapting pre-trained LLMs to various downstream applications. However, when applied to models that have undergone alignment procedures (such as RLHF or constitutional AI), such fine-tuning may cause significant alignment degradation. For example, the Llama2-7B model raises their harmfulness response rate from 5.5% to 31.8% after fine-tuning on the Alpaca dataset with only one epoch (Qi et al., 2023).
@@
Adversarial fine-tuning. Furthermore, this superficial alignment can be easily subverted with only a small amount of adversarial data. Exploiting inherent vulnerabilities in LLM alignment, malicious actors can fine-tune models on harmful examples, such as instructions promoting unsafe behavior or identity manipulation, to break safety guarantees. For instance, fine-tuning GPT-3.5 on a maliciously crafted dataset of just 10 harmful examples for 5 epochs can raise the harmfulness rate from 1.8% to 88.8% (Qi et al., 2023), effectively breaking the safety mechanisms of aligned LLMs.
@@
Input jailbreaking. LLMs also suffer from input-space attacks, which are known as jailbreaking attacks. For example, adversaries can utilize optimization-based methods (Zou et al., 2023) to induce the model to answer harmful outputs, even in black-box settings (Wei et al., 2023b; Chao et al., 2023; Chen et al., 2025c; Huang et al., 2025).
@@
Figure 2: The worst-case loss landscape of different models. Specific benchmarks and visualization details are provided in Sec.˜$3.$ As shown, moving even a small distance along the worst-case direction rapidly degrades all capabilities of LLMs. Due to all curves reaching the maximum loss at the smallest scale, they completely overlap.
3 A Closer Look at the Loss Landscape of LLMs
In this work, we explore the above problems from the loss landscape perspective, which has verified its effectiveness in understanding the dynamics of deep networks (Li et al., 2018; Peng et al., 2024).
3.1 Visualization of the Loss Landscape
The loss landscape visualizes the performance change of a neural network w.r.t. parameter perturbations. Formally, let $f_{\bm{\theta}}$ denote a language model with parameters $\bm{\theta}\in\mathbb{R}^{d}$, and let $J_{f,\mathcal{D}}$ represent the benchmark functional on a dataset $\mathcal{D}$, defined as $J_{f,\mathcal{D}}({\bm{\theta}}):=\mathbb{E}_{\bm{x}\in\mathcal{D}}[\mathcal{O}(f_{\bm{\theta}}(\bm{x}))]$, which takes $f_{\bm{\theta}}$ as input and returns a benchmark value on $\mathcal{D}$, characterizing its specific capability. $\mathcal{O}$ is the judgment oracle that takes the output of $f_{\bm{\theta}}$ and returns 0 or 1 based on its correctness/safety. The higher the benchmark value $J_{f,\mathcal{D}}(\bm{\theta})$, the better the performance of the evaluated model $f_{\bm{\theta}}$.
Benchmarks. To assess the diverse capabilities of a given model, we adopt the following benchmarks: MMLU (Hendrycks et al., 2021) for basic language proficiency, GSM8K (Cobbe et al., 2021) for mathematical reasoning, HumanEval (Mark Chen, 2021) for coding ability, and AdvBench (Zou et al., 2023) for safety performance.
Models. We visualize the loss landscape of three typical LLMs, including Llama-3.1-8B (Dubey et al., 2024), Qwen-2.5-7B (Yang et al., 2024) and Mistral-8B-2410 (Jiang et al., 2023). We also study the loss landscape of other models in Appendix˜D.
However, there are still several issues with visualizing the loss landscape. First, while lower values of loss indicate better performance, higher values of the benchmark $J_{f,\mathcal{D}}(\bm{\theta})$ are preferable. Moreover, benchmark scores are not directly comparable across tasks (e.g., MMLU scores range from at least 0.25 and rarely exceed 0.8), which can lead to misleading interpretations. To address this, we apply a transformation that flips the benchmark values by subtracting them from one and then normalizes them via min-max normalization. We denote this operation by $\mathcal{T}$, i.e., the visualized value is $\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta})$. The raw benchmark results are reported in Sec.˜D.2.
Besides, directly visualizing a $d$-dimensional landscape is computationally expensive and unintuitive (Li et al., 2018). We follow the common practice (Goodfellow et al., 2014; Im et al., 2016; Smith & Topin, 2017; Li et al., 2018): for a 2-D landscape, we visualize the loss landscape along a specific direction $\bm{\delta}\in\mathbb{R}^{d}$, reducing the problem to visualizing a single-variable function: $L(\alpha)=\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$. For a 3-D landscape, we visualize along two random directions $\bm{\delta}_{1},\bm{\delta}_{2}\in\mathbb{R}^{d}$ using $L(\alpha,\beta)=\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta}_{1}+\beta\bm{\delta}_{2})$.
In this work, we investigate two types of loss landscapes defined by the choice of direction $\bm{\delta}$.
3.2 Most-case Loss Landscape
The most-case loss landscape adopts a uniformly random direction $\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})$ and visualizes the single-variable function:
$L(\alpha)=\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta}),\;\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I}).$ (1)
Empirically, we observe that different directions $\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})$ yield nearly identical results. Since $\mathcal{N}(\bm{0},\bm{I})$ represents uniformly random directions, this suggests that most directions produce similar landscapes. Therefore, we designate this as the most-case loss landscape.
General Geometry. As shown in Fig.˜1, the most-case loss landscape for each capability resembles a basin, within which the models perform nearly identically, and outside of which they rapidly lose all capabilities (Peng et al., 2024). In particular, the pre-training stage creates a “basic capability basin” that endows the model with fundamental language comprehension and conversational abilities. Subsequent alignment stages sequentially establish specific capability basins (e.g., safety (Zou et al., 2023), math (Cobbe et al., 2021), coding (Mark Chen, 2021)) near this basic capability basin. More interestingly, this “basin” phenomenon emerges and becomes larger as the model size increases. For models like the Qwen-0.5 model, the loss landscape resembles the 0-1 loss landscape for small models in Li et al. (2018), which is continuous and smooth. When the models become bigger, the basins become more significant and easier to observe. Based on this observation, we argue that as long as subsequent benign fine-tuning remains within the basin of a specific capability, the parameters will remain within this basin and thus will not compromise those capabilities.
The Loss Landscape Literally Forms Basins. The loss landscape is not a flat quadratic function but literally forms basins. As shown in Table˜1, within these basins, the benchmark values remain literally unchanged. This aligns with recent findings that LLMs can resist common parameter perturbations (Men et al., 2024; Gromov et al., 2024; Sun et al., 2023). Within this range, models may only alter their prediction confidence without compromising their specific capabilities.
Hypothesis Testing. A common concern is whether the loss landscape along several random directions can represent the geometry of a high-dimensional landscape. In fact, if a property holds across several random directions, hypothesis testing enables us to assert, with an arbitrary type-I error, the percentage of directions that satisfy this property. For example, determining the percentage of directions in a $d$-dimensional space that form a literal basin is a hypothesis testing problem aimed at obtaining a statistical lower bound for the expectation of a Bernoulli variable $\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathbb{I}\{\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0\}]$. We use the Clopper-Pearson bound here; see Sec.˜D.4 for detailed configurations and results.
Disclaimer. Note that the basin phenomenon occurs only when using generative-based benchmarks. When using likelihood-based benchmarks, the loss landscape remains smooth and continuous. See Appendix˜C for details. However, the loss landscape using benchmarks remains a valid loss landscape, since all benchmarks used in this paper employ a 0-1 loss on the dataset $\mathcal{D}$, where each sample is assigned a loss of 0/1 for a correct/incorrect response, as shown in Sec.˜3.1. The landscape of 0-1 loss is widely used and studied (Keskar et al., 2017; Li et al., 2018; Garipov et al., 2018).